ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 231, 2019

Amaral, M. D., J. V. da Silva, G. C. Ricarte and R. Teymurazyan Sharp regularity estimates for quasilinear evolution equations

25

Bravo, M., H. Prado and E. G. Reyes Nonlinear pseudodifferential equations defined by elliptic symbols on L^{p}(ℝ^{n}) and the fractional Laplacian

269

Bruin, H., R. Fokkink and C. Kraaikamp Erratum to: “The convergence of the generalised Selmer algorithm”

505

Ciubotaru, D. and V. Heiermann On the reducibility of induced representations for classical p-adic groups and related affine Hecke algebras

379

Cook, B. Maximal function inequalities and a theorem of Birch  211
Ding, Y. Companion points and locally analytic socle for GL_{2}(L)  47

Doucha, M. Universal actions and representations of locally finite groups on metric spaces

343

Fokkink, R. See H. Bruin Galego, E. M. and A. L. Porto da Silva A solution to the

Cambern problem for finite-dimensional Hilbert spaces

419

Hanzer, M. R-groups for metaplectic groups  467

Hartmann, A., X. Massaneda and A. Nicolau Finitely generated ideals in the Nevanlinna class

139

He, J.-W. and Y. H. Zhang Local cohomology associated to the radical of a group action on a noetherian algebra

303

Heiermann, V. See D. Ciubotaru Iosevich, A. and B. Liu Equilateral triangles in subsets of ℝ^{d} of

large Hausdorff dimension

123

Klopsch, B., A. Thillaisundaram and A. Zugadi-Reizabal Hausdorff dimensions in p-adic analytic groups

1

Kraaikamp, C. See H. Bruin

Kurka, O. The isomorphism class of c_{0} is not Borel  243
Li, B., C. Schütt and E. M. Werner Floating functions  181

Liu, B. See A. Iosevich Massaneda, X. See A. Hartmann Nicolau, A. See A. Hartmann Ostaszewska, U., E. Schmeidel and M. Zdanowicz Bounded

nonoscillatory solutions of a system of dynamic equations with neutral term on time scales

489

Porto da Silva, A. L. See E. M. Galego Prado, H. See M. Bravo Reyes, E. G. See M. Bravo Ricarte, G. C. See M. D. Amaral Roychowdhury, M. K. Optimal quantization for the Cantor

distribution generated by infinite similutudes

437

Schmeidel, E. See U. Ostaszewska Schütt, C. See B. Li da Silva, J. V. See M. D. Amaral Teymurazyan, R. See M. D. Amaral Thillaisundaram, A. See B. Klopsch Werner, E. M. See B. Li Zdanowicz, M. See U. Ostaszewska Zhang, Y. H. See J.-W. He Zugadi-Reizabal, A. See B. Klopsch